OpenStudy (anonymous):

Find the amplitude and the period of b(x) = -5sin(πx).

OpenStudy (sidsiddhartha):

amplitude and period of a function like \[y=asin(bx)\] is amplitude=a period=\[2(\pi)/b\]

OpenStudy (anonymous):

okay well i'm still confused

OpenStudy (sidsiddhartha):

okay In general, if we have y = Asin(Bx - C) then, amplitude is A period is 2π/B phase shift C/B here the function is of the form y=A sin(Bx)...................(1) and ur function is y=-5sin(pi*x) compare it with equation (1) and u'll get A=-5 which is the amplitude and period=\[2\pi/B=2\pi/\pi=2\]

OpenStudy (sidsiddhartha):

ok then??

OpenStudy (anonymous):

ohh okay i see now thank you